# Genie: Smart ROS-based Caching for Connected Autonomous Robots

### v1

@@ I Introduction
Autonomous robots are rapidly expanding from a research-oriented subject to real-world applications, with autonomous vehicles being a particular and promising representation. Companies such as Uber and Waymo already started deploying fully autonomous taxi fleets in a limited number of areas [1, 2]. However, several outstanding questions remain when it comes to accuracy, latency, timing-predictability, and energy efficiency.
@@
However, the utility of edge computing for autonomous vehicles hinges on the edge provider’s ability to deliver markedly enhanced performance, particularly in terms of latency, and to provide additional, valuable information about the vehicle’s immediate environment. Moreover, it is crucial that the technology deployed by these providers is compatible with widely-used autonomous frameworks, such as the Robot Operating System (ROS) [4].
Insight: The advent of edge servers, particularly those equipped with advanced GPU capabilities, presents significant opportunities to augment computational efficiency, especially if the results of these computations can be reused effectively.
Motivated by such insight, this paper proposes a novel approach that leverages the local coherence of edge servers to address the apparent contradiction between the limitations of low-power edge servers and the demand for improved latency. Since the computations on these servers pertain to data (e.g., images) collected from the immediate vicinity, employing locality-aware caching methods could substantially reduce latency. Furthermore, the efficacy of such caching mechanisms is likely to improve over time with the accumulation of data, enabling edge providers to furnish autonomous vehicles with increasingly accurate and reliable information.
@@
Contributions. To rectify the aforementioned three challenges, we design and implement the concept of Genie: a distributed ROS-based encapsulation interface that can transparently perform object-oriented caching operations, share data with neighboring Genies, and build a three-dimensional object map of the surrounding area of the edge server cluster.
Our design of Genie has three main parts:
  • ROS-based Transparent Caching. We develop a Genie ROS node that can transparently intercept communication from/to existing ROS nodes, and perform caching operations without the need to modify the underlying software.
  • Distributed Cache Construction. We develop an algorithm that can enable the Genies located on the edge server cluster to identify and communicate with each other. Through this communication, each Genie will build and improve its local cache.
  • Object Map for Autonomous Vehicles. Finally, we identify key information types in autonomous driving and determine that real-world recognized objects in the 3D space are exceptionally useful for both removing redundant computation and providing additional information (e.g., a traffic light around the corner) for autonomous vehicles.
Implementation and Evaluation. We fully implement and evaluate our design on an edge-server cluster that replicates the vehicular edge computing platform adopted by our industry partner. This platform includes several Jetson TX2s and Jetson AGX Xaviers, in addition to relevant networking equipment. We use the prominent autonomous driving software, Autoware [5], as an open-source representative of a ROS-based end-to-end driving solution. Our evaluations show that by using Genie, the latency for the Autoware Vision Detector can be improved by 82% on average and up to 95% in a few cases. Moreover, object reusability is proven effective, offering 31% of reusable objects on average and up to 67%. By calculating a confidence score, we also measure the quality of the gathered cache and show that gathering information from multiple cars simultaneously can indeed improve the data quality in our method.
@@ II Background and Motivation
Autonomous Vehicles (AVs) rely on a multi-stage computational process that handles input from multiple sensors like cameras and LiDAR. This pipeline includes key steps—perception, localization, detection, prediction, planning, and control—that build the vehicle’s understanding of its surroundings to make safe and effective driving decisions [6]. Such computational processes must meet stringent timing constraints to ensure safety. However, state-of-the-art deep learning models introduce great computational demands even advanced hardware of edge devices still struggles to keep pace with the demanding processing needs, leading to inevitable trade-offs in data and algorithm selection to maintain safety and efficiency.
As shown in table I, we test six state-of-the-art object detection models [7, 8] in autonomous driving pipelines on three advanced GPU-enabled embedded devices (Nano, AGX, and Orin), and an edge server with an A4500 GPU. The results demonstrate that leveraging more powerful embedded devices or edge servers achieves a non-trivial speedup of up to 10.37x to meet the natural timing constraints of autonomous driving tasks. For instance, YOLOv8s [7] and DETR-ResNet-101 [8] achieve 5.02x, 10.37x speedup on the A4500 compared to its baseline on the Nano. This indicates that the advent of edge servers, particularly those equipped with advanced GPU capabilities, presents significant opportunities to augment computational efficiency, especially if the results of these computations can be reused effectively. Similar trends have been shown in prior work for other safety-critical algorithms leveraging cloud resources for AVs [9].
Intuitively, incorporating edge servers or more powerful embedded devices as distributed caches for AVs could be a promising direction to help the safety-critical autonomous driving process meet stringent timing constraints. Specifically, it can benefit by offloading computational tasks, speeding up data processing, and reducing system failures like the out-of-memory errors seen with DETR-ResNet-101-DC5 [8] on less capable Nano hardware.
TABLE I: Runtime/speedup of autonomous driving models running on different embedded devices (Nano, AGX, Orin) and edge server GPUs (A4500). OOM indicates an out-of-memory error. The slowest success execution among devices is used as the baseline to calculate speedup.
@@ III-A Design Considerations
Figure 1: An autonomous car connected to the edge.
@@
Proposed Edge Benefits. Fig. 1 shows a scenario where a car is connected to the edge. As is evident in the figure, the necessary setup for autonomous vehicles is for the vehicle itself to have every service required for full autonomy (depicted as Node X to Node D). This is a requirement because the connection to the edge-based server could be cut off at any given moment. In the scenario of Fig. 1, the network provider has created a duplicate of a service (Node Y) that already exists on the car. The red arrows in the figure indicate that every message coming out of Node X to Node Y is duplicated both on the edge and on the car. This is done by replicating all the subscribed ($T_{S}(Y)$) and published ($T_{P}(Y)$) topics of Node Y to the edge. The results are then fed into Node Z. In this configuration, it is possible for Node Z to receive duplicated messages, and has to identify and discard one of the duplicates (since the services are identical). This is the most common scenario where the edge can be useful. For example, Autoware does rely on object detection for some tracking and sign detection functionality. However, slow object detection is still serviceable without a remote connection. The value of the edge here is to run faster object detection and thus, increase the overall decision-making accuracy of Autoware. The other potentially useful scenario is for the edge to have an extra service that would improve accuracy when available, depicted as Node F in Fig. 1. For example, edge devices can offer a vision assistant for the car (as we shall discuss in Sec. IV-F). In both scenarios, the connection to the edge shall always remain optional.
@@ III-B Genie ROS Node for Non-Intrusive Caching
The main problem to address here is that of building and storing the cache. This problem is not as intuitive as it seems. Imagine Autoware, where ROS nodes are already implemented and deployed. Even though Autoware is open-source, other ROS-based autonomous driving software may not be. Thus, we would like to store the cache in a non-intrusive way without modifying the source code of Autoware’s ROS nodes. For that, we present the concept of Genie, an encapsulation made for ROS nodes, presented in Fig. 2. As is evident in the figure, a ROS node is encapsulated in a Genie interface conceptually. This is done by attaching an additional generic Genie ROS node developed by us to each ROS node except the master. This attachment procedure is done by automatically detecting the ROS topics subscribed to and published by that node (depicted as $T_{S}()$ and $T_{P}()$ in Fig. 2), switching them to topics affixed with -local so that they would only send and receive messages to the Genie node, and exposing the original topics by the Genie node itself. To other ROS nodes, this new Genie will act like the encapsulated node. However, before passing on any messages from other ROS nodes, the Genie node can check the local cache. Fig. 3 shows this design applied to the Object Detection service of an autonomous vehicle, where the service is encapsulated by Genie.
Figure 3: An example of applying Genie to object detection in an autonomous vehicle.
@@ III-C Distributed Collective Cache
Collective Cache Construction. Alg. 1 depicts how the cache is populated in detail, using the aforementioned interfaces. Whenever a message arrives on a topic, it is checked against the database (line 12). If the entry does not exist, the message is forwarded both on the -local version (line 13) and the -remote of that topic. If the entry exists in the database (line 16), the stored results are retrieved (line 17) and directly returned to the sender (line 19). Sending messages on the -remote topics constitutes broadcasting the message on the edge network if the Genie is on the edge, where other Genies with the same functionality have subscribed to the same set of topics (whereas sending message on -remote for local car-based Genies uploads the workload to the remote Genie). Upon receiving a message from neighboring Genies, the same procedure of cache miss (line 12) and cache hit (line 16) is followed. If the Genie has an entry for the request, the results will be sent back to the sender directly (line 19).
We shall explain line 9 and line 18 in our enhanced cache design, discussed in Sec. III-D.
@@
10:       BuildHashMap(T) $\triangleright$ Create a new hashmap for never-before-seen topics.
@@ III-D Driving-specific Caching
6:    if type_of(M) $\in TYPES$ then $\triangleright$ Recognized type.
@@
10$\triangleright$ Boost confidence for existing objects.
11:          else
12:             OBJECT_MAP[M.data.location].add(object)
13:function EnhancedCacheBoostData(cacheEntry)
14:    if type_of(M) $\in$ TYPES then $\triangleright$ Recognized type.
15:       for object $\in$ cacheEntry.value.objects do
16:          for stored_object $\in$ OBJECT_MAP[object.location] do
@@
18:                cacheEntry.value.objects.add(stored_object)
19$\triangleright$ Add all the new objects we know about at this location.
@@ IV Evaluation
IV-A Setup
Devices. In this section, we would like to evaluate the efficacy of our design in terms of tail latency, reusability, and value-added accuracy. For that purpose, we use an edge server cluster that replicates the platform adopted by our industry partner. This cluster consists of three NVIDIA AGX Xaviers and three NVIDIA Jetson TX2s. We use the Jetson TX2s to represent cars, meaning that they will run the full autonomous driving software plus the ROS master. We use AGX Xaviers as remote edge servers. For heterogeneous configurations, we also use a powerful server with an Intel Xeon E5-2650 CPU and an NVIDIA Quadro RTX 4000.
Configuration. We specifically co-located our cars and edge servers to eliminate network delays. This is because our methodology is not designed to improve network delays. Instead, we would like to purely measure the effect of our caching method on computational latency. We use Autoware as the representative autonomous driving software. Autoware was chosen because it is open-source and is gaining considerable traction in the industry [12]. However, other ROS-based autonomous driving software also exists from companies such as BMW, Baidu, etc [13, 14] that have the same basic architecture as Autoware.
Data. We use the KITTI Vision Benchmark suite [15] raw data to playback recorded data from onboard sensors of vehicles. This data includes four camera streams, a LIDAR point cloud stream, and precise location data.
Measurements. We measure tale-latency in terms of milliseconds (ms), image reusability in terms of the ratio between cache hits and total cache requests on the image cache, object reusability in terms of cache hits to total cache requests on the object map, and confidence boost in terms of the average sum of total boost for the object map.
Scenarios. To cover various scenarios of autonomous driving, we evaluate three complexities: 1 car, 2 cars, and 3 cars. The more cars that are involved and communicating, the more information can our collective cache gather.
Homogeneous vs. Heterogeneous Cluster. Although we strongly believe that next-generation edge server clusters should only contain low-power devices, we wanted to test the effect of adding a powerful server to the cluster. In our tests marked as heterogeneous throughout this section, we use our Quadro-based server as one of the remote machines (e.g., in the 1-car scenario, the remote Genie and the corresponding remote Autoware Vision Detector are executed on the powerful server).
Case Study of Genie. Finally, we conduct a case study on how Genie can be useful in vision-assisted cars by enabling communication with vision-enabled remote Genies.
@@ IV-B Tail-Latency Performance
(ii) Heterogeneous setting.
Fig. 5f shows the cumulative distribution function of response time for Genie (DG) compared to ROS remote execution (R) and local execution (L) of Autoware Vision Detector (AVD) (also known as Autoware object detector) for the 1 car, 2 cars, and 3 cars scenario where the edge cluster is homogeneous and heterogeneous. The practical 33ms deadline is also labeled with a dashed line. As is evident in the figure, Genie outperforms both the remote execution and local execution of the object detector in all cases by 82% on average. Moreover, Genie is mostly capable of meeting real-time demand even with the limited hardware at hand. Furthermore, Genie is much more consistent in terms of response time than purely running the AVD, since cache hit and miss and executing both Alg. 1 and Alg. 2 are simple operations and AVD is more selectively executed under DG compared to R and L. Moreover, the addition of cars does not incur noticeable overhead, even though the latency for network communication is also included in the measurements (albeit on wired Ethernet networks).
In the case of the heterogeneous server cluster under the 1-car scenario, the Quadro RTX 4000 remote server (marked with R) is capable of executing the vision tasks with a runtime similar to Genie. However, as soon as additional cars are added, the ad-hoc nature of the remote execution leads to substantially worsened average and worst-case latency. Another key observation is that the consistency of Genie execution time is apparent between the low-power embedded servers in the homogeneous scenario and the powerful server in the heterogeneous scenario. This consistency (which is because CPU operations required by Genie are short and bursty, and both ARM and Intel CPUs are generally capable of adequately executing them in a similar time [16]) can give a good assurance to system developers that might consider using our method on their desired architecture.
@@ IV-D Object Caching
As we discussed in Sec. III-D, object mapping can be more useful for autonomous driving. The drawback is that this kind of cache is not as general and scalable as our message-based cache. Fig. 7 shows the average object reusability (i.e., cache hits on objects, versus total number of requests). As is evident in the figure, the reusability rate is substantially higher (as much as 4 times) compared to the image cache, reaching 31% on average and a maximum of 67%.
Interestingly, in the homogeneous case, the trend for reusability is the opposite of the trend in image cache. In the case of object reusability, the more cars that are present, the less reusable data is seemingly available. However, this phenomenon is since each location can contain a proportionally high number of objects. The more cars that are available means that substantially more objects will be available. However, not all of them will pass the high confidence requirement mentioned in Sec. III-D. Thus, the overall reusability slightly falls. However, the quality of the data is substantially improved as a result.
Another factor is that for the 1-car scenario, the local object map has a better reusability ratio. This is due to the selective uploading method, where not all objects are requested from the remote server (if they are cached locally).
Finally, as is evident, the very powerful server in the heterogeneous scenario has vastly enhanced the object reusability, especially when more cars are available. Moreover, the trend is as expected, where more cars lead to more reusability because the server is capable of processing 4 times the images (and thus 4 times the objects) compared to the Jetson AGX Xavier (which outweighs the drawback observed in the homogeneous architecture).
@@ IV-E Confidence Boost
We also measure how our confidence-boosting method is affected by the various scenarios. The results are recorded in Fig. 8 as the cumulative average of the score in addition to object arrays of outgoing images. Fig. 8 compares the 1 car, 2 cars, and 3 cars scenarios under both the homogeneous and the heterogeneous architectures. The values for confidence boost are measured both from remote Genies (R) and local Genies (L) for each scenario. The y-axis value is a running average of the confidence addition for each object and the x-axis is the number of recorded events (e.g., received requests). As is evident in the figure, local Genies have generally less confidence because their communication is limited to the remote Genie. Moreover, the 3-car scenario naturally features a slightly higher confidence boost compared to the 2-car scenario. The same is true for the 2 cars scenario versus the 1-car scenario. Over a long period, the more cars that are present in the system, the higher the quality of the data on the remote Genies will become. Moreover, the heterogeneous architecture can achieve a slightly higher cumulative average value due to the faster processing of the powerful server.

### v2

@@ I Introduction
Autonomous robots are rapidly expanding from a research-oriented subject to real-world applications, with autonomous vehicles being a particular and promising representation [1, 2, 3, 4]. Companies such as Waymo already started deploying fully autonomous taxi fleets in a limited number of areas [3, 4]. However, several outstanding questions remain when it comes to accuracy, latency, timing-predictability, and energy efficiency.
@@
Key Insight: Utilizing edge servers with locality-aware caching can reduce latency significantly if results can be reused effectively, making it feasible to deliver enhanced performance and provide additional, valuable environmental information to autonomous vehicles.
@@
Contributions: To address these issues, we introduce Genie, a distributed ROS-based interface for object-oriented caching and data sharing, constructing a 3D object map of the edge server’s surroundings: (1) ROS-based Transparent Caching: Genie intercepts ROS communications for caching without modifying software. (2) Distributed Cache Construction: An algorithm allows Genies to communicate, enhancing their local caches. (3) 3D Object Map: Identifies and caches useful real-world objects for autonomous vehicles, reducing redundant computation and providing extra information.
Implementation and Evaluation: We implemented Genie on a vehicular edge computing platform with Jetson TX2s and AGX Xaviers, using Autoware [6] as a representative ROS-based system. Our evaluation shows that Genie reduces latency by 82% on average, with a peak improvement of 95%. Object reusability is effective, with 31% reusable objects on average and up to 67%. A confidence score demonstrates improved data quality through multi-car information gathering.
@@ II Background and Motivation
Autonomous Vehicles (AVs) rely on a multi-stage computational process involving sensors like cameras and LiDAR to ensure safe driving decisions [7]. This pipeline, which includes perception, localization, detection, prediction, planning, and control, must meet stringent timing constraints to maintain safety. However, state-of-the-art deep learning models impose heavy computational demands, and even advanced edge devices struggle to keep pace, forcing trade-offs in data and algorithm selection.
Our evaluation of six cutting-edge object detection models [8, 9] across various GPU-enabled devices (Nano, AGX, Orin) and an edge server with an A4500 GPU shows significant speedups. For instance, YOLOv8s and DETR-ResNet-101 achieve speedups of 5.02x and 10.37x on the A4500 compared to the Nano, illustrating that advanced GPU-equipped edge servers can substantially enhance computational efficiency by reusing results. Note that for some small models like YOLOv8s, even all evaluated devices provides acceptable latency, but this model has the lowest accuracy.
The advent of edge servers presents a significant opportunity to augment computational efficiency. By offloading tasks to edge servers, AVs can speed up data processing and avoid system failures, such as out-of-memory errors observed with DETR-ResNet-101-DC5 on less capable hardware.
This motivates the incorporation of edge servers or powerful embedded devices as distributed caches to help AVs meet stringent timing constraints, enhance safety-critical processes, and improve overall performance.
TABLE I: Runtime/speedup of autonomous driving models on different devices (Nano, AGX, Orin) and edge server GPU (A4500). OOM indicates an out-of-memory error. The baseline for speedup is the slowest successful execution among devices.
@@ III-A Design Considerations
Proposed Edge Benefits. Fig. 1 shows a scenario where a car is connected to the edge. As is evident in the figure, the necessary setup for autonomous vehicles is for the vehicle itself to have every service required for full autonomy (depicted as Node X to Node D). This is a requirement because the connection to the edge-based server could be cut off at any given moment. In the scenario of Fig. 1, the network provider has created a duplicate of a service (Node Y) that already exists on the car. The red arrows in the figure indicate that every message coming out of Node X to Node Y is duplicated both on the edge and on the car. This is done by replicating all the subscribed ($T_{S}(Y)$) and published ($T_{P}(Y)$) topics of Node Y to the edge. The results are then fed into Node Z. In this configuration, it is possible for Node Z to receive duplicated messages, and has to identify and discard one of the duplicates (since the services are identical). This is the most common scenario where the edge can be useful. For example, Autoware relies on object detection for tracking and sign detection functionality. However, slow object detection is still serviceable without a remote connection. The value of the edge here is to run faster object detection and thus, increase the overall decision-making accuracy of Autoware. The other potentially useful scenario is for the edge to have an extra service that would improve accuracy when available, depicted as Node F in Fig. 1. For example, edge devices can offer a vision assistant for the car (as we shall discuss in Sec. IV-F). In both scenarios, the connection to the edge shall remain optional.
Figure 1: Left: an autonomous car connected to the edge. Right: an example of applying Genie in an autonomous vehicle.
@@ III-B Genie ROS Node for Non-Intrusive Caching
The main problem to address here is that of building and storing the cache. This problem is not as intuitive as it seems. Imagine Autoware, where ROS nodes are already implemented and deployed. Even though Autoware is open-source, other ROS-based autonomous driving software may not be. Thus, we would like to store the cache in a non-intrusive way without modifying the source code of Autoware’s ROS nodes. For that, we present the concept of Genie, an encapsulation made for ROS nodes, presented in the right of Fig. 1. As is evident in the figure, a ROS node is encapsulated in a Genie interface conceptually. This is done by attaching an additional generic Genie ROS node developed by us to each ROS node except the master. This attachment procedure is done by automatically detecting the ROS topics subscribed to and published by that node (depicted as $T_{S}()$ and $T_{P}()$ in Fig. 2), switching them to topics affixed with -local so that they would only send and receive messages to the Genie node, and exposing the original topics by the Genie node itself. To other ROS nodes, this new Genie will act like the encapsulated node. However, before passing on any messages from other ROS nodes, the Genie node can check the local cache. Fig. 1 shows this design applied to the Object Detection service of an autonomous vehicle.
@@ III-C Distributed Collective Cache
Collective Cache Construction. Alg. 1 depicts how the cache is populated in detail, using the aforementioned interfaces. Whenever a message arrives on a topic, it is checked against the database (line 12). If the entry does not exist, the message is forwarded both on the -local version (line 13) and the -remote of that topic. If the entry exists in the database (line 16), the stored results are retrieved (line 17) and directly returned to the sender (line 19). Sending messages on the -remote topics constitutes broadcasting the message on the edge network if the Genie is on the edge, where other Genies with the same functionality have subscribed to the same set of topics (whereas sending message on -remote for local car-based Genies uploads the workload to the remote Genie). Upon receiving a message from neighboring Genies, the same procedure of cache miss (line 12) and cache hit (line 16) is followed. If the Genie has an entry for the request, the results will be sent back to the sender directly (line 19). We shall explain line 9 and line 18 in our enhanced cache design, discussed in Sec. III-D.
@@
10:       BuildHashMap(T) $\triangleright$ Create a new hashmap for never-seen topics.
@@ III-D Driving-specific Caching
6:    if type_of(M) $\in TYPES$ then
@@
10:          else
11:             OBJECT_MAP[M.data.location].add(object)
12:function EnhancedCacheBoostData(cacheEntry)
13:    if type_of(M) $\in$ TYPES then
14:       for object $\in$ cacheEntry.value.objects do
15:          for stored_object $\in$ OBJECT_MAP[object.location] do
@@
17:                cacheEntry.value.objects.add(stored_object)
@@ IV Evaluation
IV-A Experimental Setup
Devices. We evaluate our design’s efficacy in tail latency, reusability, and value-added accuracy using an edge server cluster modeled after our industry partner’s platform. The cluster includes three NVIDIA AGX Xaviers and three NVIDIA Jetson TX2s. The Jetson TX2s represent cars running the full autonomous driving software plus the ROS master, while the AGX Xaviers serve as remote edge servers. For heterogeneous setups, we add a server with an Intel Xeon E5-2650 CPU and NVIDIA Quadro RTX 4000.
Configuration. Cars and edge servers are co-located to eliminate network delays, focusing solely on computational latency effects from our caching method. We use Autoware, an open-source autonomous driving software gaining industry traction [13]. Other ROS-based alternatives like those from BMW and Baidu share similar architectures [14, 15].
Data. The KITTI Vision Benchmark suite [16] provides raw data for playback from vehicle sensors, including four camera streams, a LIDAR point cloud, and precise location.
Measurements. Metrics include tail latency, image reusability (cache hits/total cache requests for image cache), object reusability (cache hits/total cache requests for object map), and confidence boost (average total boost for object map).
Scenarios. We evaluate scenarios with 1, 2, and 3 cars to explore different levels of complexity. Increased car involvement enhances collective cache information.
Homogeneous vs. Heterogeneous Cluster. We test clusters with low-power devices and add a powerful server for heterogeneous configurations. The Quadro-based server acts as one of the remote machines in heterogeneous tests.
Genie Case Study. We perform a case study demonstrating how Genie aids vision-assisted cars through communication with vision-enabled remote Genies.
@@ IV-B Tail-Latency Performance
Fig. 4 shows the cumulative distribution function (CDF) of response times for Genie (DG) versus ROS remote (R) and local (L) execution of the Autoware Vision Detector (AVD) across scenarios with 1, 2, and 3 cars in homogeneous and heterogeneous edge clusters. The 33ms deadline is marked with a dashed line. Genie outperforms both remote and local executions of AVD, achieving an average improvement of 82%, consistently meeting real-time demands with stable response times due to efficient caching and selective execution of algorithms. Adding more cars does not significantly impact latency, even with network communication, due to the wired Ethernet setup.
In the 1-car heterogeneous scenario, the server (R) performs similarly to Genie initially, but as cars increase, remote execution experiences higher latency, showcasing Genie’s advantage in maintaining consistent performance. Genie’s execution time remains stable across low-power embedded and powerful heterogeneous servers, thanks to its efficient CPU operations, adaptable to both ARM and Intel CPUs [17]. This ensures Genie’s reliability across various architectures.
@@ IV-D Object Caching
As discussed in Sec. III-D, object mapping is particularly beneficial for autonomous driving, despite being less general and scalable than our message-based cache. Fig. 6 illustrates the average object reusability, defined as the ratio of cache hits on objects to total requests. The reusability rate for object caching is significantly higher than for image caching, averaging 31% and reaching a maximum of 67%, which is four times greater than image cache rates. In homogeneous scenarios, object reusability decreases with more cars because each location can contain a large number of objects. While more cars increase the number of available objects, not all meet the high confidence threshold from Sec. III-D, slightly reducing overall reusability but enhancing data quality. For the 1-car scenario, the local object map shows better reusability due to selective uploads that avoid unnecessary requests to the remote server when objects are cached locally. In heterogeneous scenarios, the powerful server significantly enhances object reusability, especially with more cars. The server can process four times as many images (and thus objects) compared to the Jetson AGX Xavier, overcoming the limitations seen in the homogeneous setup and leading to increased reusability with additional vehicles.
@@ IV-E Confidence Boost
We measured the confidence boost of our method across various scenarios, as shown in Fig. 7. The cumulative average scores results are compared for 1, 2, and 3 cars in both homogeneous and heterogeneous architectures. Confidence boosts are recorded for both remote (R) and local (L) Genies. The y-axis shows the running average confidence addition per object, while the x-axis represents the number of recorded events. Local Genies have generally lower confidence due to limited communication. The 3-car scenario shows a higher confidence boost compared to 2-car and 1-car scenarios, and the heterogeneous architecture achieves a higher cumulative average due to faster processing by the powerful server. Over a long period, the more cars that are present in the system, the higher the quality of the data on the remote Genies will become. Moreover, the heterogeneous architecture can achieve a slightly higher cumulative average value due to the faster processing of the powerful server.
(a) Homogeneous setting.
(b) Heterogeneous setting.